The width of a rectangle is 13 feet less than the length. If the area is 30 square feet, find the width.

Answer by TimothyLamb(4379)

---
y = x - 13
xy = 30
---
xy = 30
x(x - 13) = 30
xx - 13x - 30 = 0
---
the above quadratic equation is in standard form, with a=1, b=-13, and c=-30
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 -13 -30
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = 15
x = -2
---
the negative root doesn't fit the problem statement, so use the positive root:
---
x = length = 15 ft
y = width = 2 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php